Zip Games purchases blank DVD disks onto which it copies its software for sale through its mail order operation. The firm opens every day. A disk costs Zip $0.20. Processing an order for more disks cost $15. Zip uses 60000 disks annually, and the company has a 25% holding cost rate. Show steps to find:

a. EOQ.

b. What is the reorder point if lead time is 2 days?

c. What is the cycle time?

d. What is the total inventory cost?
Answer #1

Question: Zip Games purchases blank DVD disks onto which it copies its software for sale through its mail order operation. The firm opens every day. A disk costs Zip $0.20. Processing an order for more disks cost $15. Zip uses 60000 disks annually, and the company has a 25% holding cost rate. Show steps to find:

Answer:

a. EOQ. = 6000 Units

b. What is the reorder point if lead time is 2 days = 328.76 Units - [329 Units (rounded)]

c. What is the cycle time? = 36.5 Days - [37 Days (rounded)]

d. What is the total inventory cost? = $750

Explanation:

It is given that:

Demand (D) = 60000
Order Cost (S) = $15
Holding Cost (I) = 25%
Unit Cost (C) = $0.20
Lead Time (LT) = 2 days
Number of days per year = 365 (Since firm open every day)

a. EOQ.

Economic Order Quantity (EOQ) = ( 2 × D × S / H ) 1/2

Here,

D = 60000

S = 15

H = (I x C) = (25% x 0.20) = 0.05

Hence,

EOQ = Q = (2 x 60000 x 15 / 0.05)1/2

EOQ = Q = 6000 Units

b. What is the reorder point if lead time is 2 days?

Reorder Point (ROP) tells when to order.

ROP = (Demand per day (d)) x (Lead time(Lt))

Here,

Lt = Lead Time = 2 days

Demand per day (d) = Demand per year (D) / Number of working days

d = D / No of working days

d = 60000 / 365

d = 164.38

Hence,

ROP = d x Lt = 164.38 x 2

ROP = 328.76 Units

c. What is the cycle time?

Cycle Time = (Number of working days x Q) / D

Here,

Number of working days = 365

EOQ = Q = 6000

D = 60000

Hence,

Cycle Time = (365 x 6000) / 60000 = 2190000 / 60000

Cycle Time = 36.5 Days

d. What is the total inventory cost?

Total Inventory Cost = TIC = C (Q / 2) + S (D / Q)

Here,

C = 0.20

Q = 6000

S = 15

D = 60000

Hence,

TIC = 0.20 (6000 / 2) + 15 (60000 / 6000)

TIC = 600 + 150 = 750

Total Inventory Cost = $750